OpenStudy (anonymous):

Find the perimeter of xyz x(-6,2) y(2,8) x(-6,8) I forgot the distance thingy formula

OpenStudy (lgbasallote):

\[\text{distance} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (lgbasallote):

unless it's 3d...

OpenStudy (lgbasallote):

it becomes \[\sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2 + (z_2 - z_1)^2}\]

OpenStudy (anonymous):

it's not 3-d haha okay so now can you help me with the question? ahah OMGOOMGOMOGMGOOGMG I LOVVEEE AVATARRRRRRRRRRRRRRR I'M SO OBSESSEEDDD LOL

OpenStudy (lgbasallote):

lol good for you :) and i really cant help because i cannot imagine what your question is

OpenStudy (anonymous):

oh trianglexyz

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

hello guys?

OpenStudy (lgbasallote):

seriously...draw it...

OpenStudy (anonymous):

they don't show a drawing. they literilly just say find the perimeter of xyz.

OpenStudy (anonymous):

If I will give you the formula then can you do it ??

OpenStudy (anonymous):

You know the formula very much all just you have to do is to use that 3 times. Can you do that @Steeben

OpenStudy (anonymous):

so it's rad(2+6)^2+(8-2)^2

OpenStudy (anonymous):

64+36 = 100 then 10?

OpenStudy (anonymous):

Go ahead use this for other 2 points left..

OpenStudy (anonymous):

rad(-6-2)^2+(8-8)^2

OpenStudy (anonymous):

64... 8

OpenStudy (anonymous):

This one is one distance you have found..